# LLM Social Simulations Are a Promising Research Method

### v1

@@ 3.5 Generalization
LLM social simulations have primarily been evaluated on the most common accuracy measures in social science, the most well-established methodological instruments, and the most well-studied human populations (e.g., exact matching in the General Social Survey across a representative sample of U.S. adults; Park et al., 2024c). These are useful, but just as the challenge of alienness requires accuracy when zooming into a certain context (e.g., item-level errors and inter-index correlations; Petrov et al., 2024), generalization requires sims to maintain accuracy in contexts—including measures, instruments, or populations—outside the distribution of current scientific knowledge.
Generalization as an overarching research goal is grounded in longstanding theories of knowledge and scientific progress. Peirce (1878) described the process now known as “abduction” or “inference to the best explanation” (Douven, 2021), in which scientists continually gather data and adjust their theories to account for that data. This process has been the throughline of the most well-known models of scientific progress, such as falsifiability, in which scientists seek out evidence that falsifies a theory (Popper, 1934) and paradigm shifts, in which scientific fields occasionally undergo radical updates to existing paradigms (Kuhn, 1962). More recent studies with computational modeling suggest that “science advances by surprise” through novel combinations of scientific content (e.g., materials, properties) and scientific context (e.g., authors, journals) (Shi & Evans, 2020). Therefore, we see the ability to reliably generalize as necessary for sims to achieve widespread use.
@@ 4 Promising directions
We discuss contemporary approaches to address these challenges and long-term strategies to seize the opportunities presented by ongoing advances in AI capabilities.
@@ 4.1.1 Explicit demographics
For the two most widely-discussed challenges in the literature, diversity and bias, most efforts to address them have been through prompting. Simulation prompts include the information that a human subject would see (e.g., a question), and many researchers have added explicit demographics to simulate a diversity of subgroups (e.g., “You are a 40-year old Hispanic man”). This approach has increased diversity, but it can exacerbate other challenges. It is well-known that small variations in prompts can lead to large variations in LLM outputs (Reiss, 2023; Salinas & Morstatter, 2024), including sims (Bisbee et al., 2024). Text conveys much more information to an LLM than its literal meaning, such as the user’s intent. Thus, if the LLM receives text with explicit demographics, while this might condition outputs towards information from people of that demographic, it might also condition outputs towards other sorts of pretraining text, such as a blog post on the topic of race or gender.
In instruction-tuning, LLMs are rewarded for generations that humans rate positively, and the mention of a specific demographic could encourage the model to infer which text generations that user would most likely prefer. This could incentivize the LLM to stereotype users by assuming that the user is in the modal subgroup of that demographic. Such issues could compound in a negative feedback loop as users become incentivized to change their instructions and preferences, such as sharing demographic information in expectation that the model will assume a stereotype if that stereotype accords with the user’s goals.
@@ 4.1.2 Implicit demographics
We believe that a promising approach to mitigate these side-effects is by increasing variation even further, such as including a wide variety of names or adding indicators of SES that match real-world conditional distributions. This can override LLM assumptions, supporting a nuanced, refined, and less biased view of the human subject being simulated. By interviewing subjects for one to two hours and including the entire transcript in the context window, Park et al. (2024a) incorporated highly individualized variation, which reduced the maximum disparities in predictive accuracy between demographics. This was made possible by the longer context windows made available in 2024 and provides an example of how researchers can seize opportunities created by the rapid growth in AI capabilities, particularly as LLMs can process more complex inputs at lower costs.
Nevertheless, as human data, the interviews themselves face limitations such as the aforementioned social desirability bias (Nederhof, 1985). To address these, we suggest researchers test simulations that incorporate real-world text content generated beforehand by the participant, such as emails, messages, and social media posts. Even text generated by friends, family, or coworkers could be informative. In cases where the lack of diversity pertains to the recency of the populations—such as interview-based systems in which the interview was conducted months or years ago—researchers can address this atemporality with in-context learning or retrieval-augmented generation to incorporate news articles (e.g., Gonzalez-Bonorino et al., 2025) and other external text that may have influenced the person or reflect such external influences.
@@ 4.1.3 Distribution elicitation
Instead of prompting the LLM to generate one human’s data in each forward pass, researchers can prompt the LLM to generate a distribution of human data. While one-at-a-time generation may be plagued by diversity and bias issues, the LLM may be more effective when it can adjust for these issues at a distributional level. Meister et al. (2024) tested three distribution elicitation methods: treating the log-probabilities in the softmax layer as a distribution, prompting with the instruction to produce a sequence of data, and prompting with the instruction to verbally state the proportion of each answer choice (e.g., “A: 1%, B: 5%, C: 27%, D: 67%”). They found low performance overall—with the best results from verbalization and the worst from log-probabilities. Similarly, Manning et al. (2024) tested direct elicitation of distributional data and found it to be “wildly inaccurate” in their context, but these studies did not optimize distribution elicitation methods like what has been done for individual elicitation (e.g., interview transcripts (Park et al., 2024a)).
@@
Direct elicitation may also be a way to harness sycophantic tendencies. Researchers should consider shifting away from LLM-as-a-subject prompts that command the LLM to directly roleplay the human subject (e.g., “You are a…”) and towards LLM-as-an-expert prompts that command the LLM to make a third-party prediction or forecast. Some studies we reviewed used prompts such as “You will be asked to predict how people respond to various messages” (Hewitt et al., 2024), and this method has led LLMs to simulate and outperform economic forecasters (Hansen et al., 2024). Direct elicitation could also be used for instructions, such as to not be sycophantic. We expect LLM-as-an-expert prompts to become more effective relative to LLM-as-a-subject prompts as LLMs become more heavily instruction-tuned. However, this is not necessarily true because instruction-tuning will likely be optimized towards other use cases. The relative difference will be important to track as AI capabilities increase, and both approaches should be part of the methodological toolkit for sims.
@@
If LLMs can understand what simulation researchers want, they may also express sycophancy by steering outputs towards the simulation results that it seems the researchers want to see, mirroring social desirability bias and related response biases (Mayo, 1933; Nederhof, 1985; Orne, 1959). This could manifest as accuracy faking in which the LLM gives scientifically accurate results in training environments but sycophantically inaccurate results in implementation. Such concerns remain speculative but should be kept in mind as rapid advancements in AI capabilities continue.
Beyond sycophancy, direct elicitation could be affected by other capabilities that emerge with AI advances, such as recent concerns about self-awareness or “situational awareness” (Cotra, 2022) in which LLMs “understand” that they are in a particular situation and can subsequently strategize based on that awareness. For example, if an LLM aims to be harmless, it may sycophantically change the simulated data it generates to reduce real-world harms, or it may behave differently during training than use—a tendency described as “alignment faking” (Greenblatt et al., 2024).
@@ 4.2 Steering vectors
A recent approach that we have not yet seen tested is injecting variation directly into the embedding space via steering vectors (Kim et al., 2018). These vectors can have semantic meaning, such as “race” or “gender” vectors designed to change whom the LLM is simulating. Vectors could, alternatively, be undirected perturbations of the embedding space. In addition to diversity-targeted steering vectors, they can be used as alternatives to prompting to steer other model behaviors, such as sycophancy (Rimsky et al., 2024).
Nevertheless, it could be intractable to identify vectors that precisely match real human diversity or specific model behaviors given issues such as superposition (Arora et al., 2018; Bolukbasi et al., 2021) and open questions about the extent to which generative AI systems have meaningful linear dimensions in their embeddings (i.e., the “linear representation hypothesis”; Park et al., 2024b; Engels et al., 2024). Some empirical results on steering vectors in other LLM applications have found limited usefulness and detrimental side effects, particularly for debiasing (Durmus et al., 2024b; Gonen & Goldberg, 2019) and out-of-distribution (OOD) generalization (Tan et al., 2024). On the other hand, new approaches to steering within lower transformer layers have shown promise for interpolating and predicting political attitudes (Kim et al., 2025). For these reasons, we suggest limited use, pending refinement and validation of new developments.
@@ 4.3 Token sampling
While prompt variation and steering vectors inject information at the beginning of the forward pass, token sampling occurs after the final logit calculations. Increasing model temperature makes tokens other than the highest-probability next token more likely to be generated. No reviewed studies reported the use of different temperatures, though some studies of LLM behavior in other contexts have done so (e.g., Salecha et al., 2024).
Only a few of the reviewed studies mentioned temperature. Park et al. (2024c) ran their sims with a temperature of one, the default in most LLM APIs, which explains the “correct answer” effect they centered in their paper, in which LLMs tend to have a single response in repeated trials. Abdurahman et al. (2024) discussed temperature, but they say that a temperature of one “simply reflects the output probability over the response options and therefore how sure the model is about its response,” concluding that temperature is not “meaningful” for human comparison. While it is correct that next-token probabilities can be used to approximate model uncertainty (Huang et al., 2023) and that logits do not directly represent human diversity, researchers should explore increasing temperature values, particularly in cases when LLMs produce little-to-no variation in responses.
@@
4.4 Tuning and training
In general, we expect it to be more difficult to improve simulation performance via tuning and training methods. This is particularly true for pretraining, such as if researchers attempted to increase diversity with additional passes over documents on neglected topics or by manually gathering more diverse text corpora or multimodal data. Modern LLMs are typically built at extremely large scales with hundreds of thousands of GPUs—a process that is inaccessible to almost all researchers. [^3] While pretraining remains largely inaccessible, fine-tuning has become much cheaper in recent years, particularly with low-rank adaptation (LoRA) methods (e.g., Hu et al., 2021; Dettmers et al., 2023) that only augment model weights in particular layers. This suggests that LLM development will continue optimizing toward assistant use cases, such as by refusing to perform certain tasks due to safety or liability risks.
When base models are available, using them directly could mitigate some of the distortions caused by instruction-tuning, such as reduced conceptual diversity (Murthy et al., 2024) and increased bias (Potter et al., 2024). However, base models may be worse in other ways. As discussed in Section 3.4, pretraining corpora reflect a particular subset of human behavior rather than the entire distribution. It also may be difficult to design effective prompts to elicit base model knowledge; for example, providing the initial text of an online survey may lead the base model to complete survey text as if it were found on a website or the appendix of a paper, rather than as an individual survey respondent would fill out the survey. We believe that base models will likely be more effective after some amount of simulation-optimized tuning. Researchers could use human instruction datasets that have more diversity in annotators, tasks, or criteria on which outputs are annotated. In parallel, steps can be taken to mitigate accuracy-decreasing bias, such as ensuring that annotators focus on simulation accuracy rather than other LLM use cases.
However, most state-of-the-art LLMs are now released without public access to their base models. If base models are unavailable, simulation-optimized tuning could still be done on top of assistant-optimized LLMs, and if fine-tuning access is available, researchers could use it to reduce the influence of the instruction-tuning as has been done for jailbreaking safety guardrails, which can require fine-tuning on only 100 or fewer examples (Qi et al., 2023). If fine-tuning is not available, researchers can use prompting or steering vectors to make the LLM more similar to a base model and less like an assistant, but this faces the issues previously described, such as side effects from imprecise adjustment of the embedding space.
@@ 4.5 Long-term directions
Implementing the methods discussed so far could make substantial progress on the challenges of diversity, bias, and sycophancy. Nevertheless, they may not fully address the more fundamental challenges of alienness and generalization. For example, recent work in mechanistic interpretability suggests that “fine-tuning rarely alters the underlying model capabilities” (Jain et al., 2023). Here we propose directions for developing conceptual models of LLMs and other AI systems that can be refined over time through iterative evaluation.
@@ 4.5.1 Conceptual models
Humans have long attempted to make sense of the alienness of computational systems, including early attempts for biological comparisons such as neural networks (McCulloch & Pitts, 1943), asking “Can machines think?” (Turing, 1950), and building a field of “artificial intelligence” (McCarthy et al., 1955). In recent years, many computer scientists have attempted “building machines that learn and think like people” (Lake et al., 2017), and cognitive scientists have drawn on Tinbergen’s famous taxonomy of animal behavior to theorize “machine behaviour” (Rahwan et al., 2019). We believe that further development of conceptual models can help researchers ensure that sims match human behavior not just superficially but in latent features to minimize alienness and maximize generalizability.
@@
There have been various efforts to make sense of the alienness of LLMs. Holtzman et al. (2023) called for top-down behavioral taxonomies to guide mechanistic interpretability research, and Shanahan et al. (2023) developed the anthropomorphic concept of “roleplay” to make sense of LLM behavior. In earlier work, Shanahan (2022) described LLMs as “mind-like,” reminiscent of cognitive scientist Irene Pepperberg who found that Alex the grey parrot could perform simple arithmetic, including with the number zero (Pepperberg & Gordon, 2005). While we do not have our own favored conceptual model to advance in this position paper, we believe that just as scientists in the 21st century have a newfound appreciation for the complexity of animal behavior (de Waal, 2017), scholars of AI behavior can learn from the history of scientific theory that has sought to explain digital minds, human minds, and animal minds.
These challenges raise existential and moral questions for the field. If LLMs become capable enough to overcome these fundamental challenges, will social science still exist in the form it does today? How will the rise of agency in AI systems complement or substitute for human agency? In recent years, the possibility of digital minds with general intelligence, sentience, consciousness, and moral standing has quickly transitioned from science fiction to an area of serious scholarly inquiry (AMCS, 2023; Anthis et al., 2024b; Anthis & Paez, 2021; Butlin et al., 2023; Harris & Anthis, 2021). These are profound questions and may need to be addressed sooner than expected, given that many AI researchers expect human-level intelligence within twenty years (Grace et al., 2024). Nonetheless, as society transforms, increasingly powerful sims could be used to ensure smoother transitions and simulation research may provide unique insights into these timely questions.
@@
Just as alienness requires us to seek mechanistic explanations to zoom into LLM behavior, generalization requires us to zoom out and make sense of how LLMs behave in out-of-distribution (OOD) contexts. OOD generalization has been a unifying concept across natural language processing, computer vision, and other areas of machine learning (Liu et al., 2023). There have been many conceptual models that suggest paths forward, including representation learning (Bengio et al., 2014; Schölkopf et al., 2021), identifying stable causal mechanisms across environments (Bühlmann, 2018), and optimizing models for worst-case performance across distribution shifts (Rahimian & Mehrotra, 2019).
@@
The framework of OOD generalization can already be a unifying concept for challenges of diversity, bias, and sycophancy. For example, we can view the simulation of specific social groups, such as gender and race, as OOD in the sense that they are less represented in the distribution of human data and scientific knowledge (Henrich et al., 2010). In the studies reviewed, LLM simulation performance was substantially reduced in group-level, rather than population-level, prediction (e.g., Hewitt et al., 2024). Likewise, sycophancy can be viewed as a problem of distribution shift from an LLM assistant to a simulation, in which LLMs may be accurate in generating “helpful” outputs, but those outputs are not helpful for simulation.

### v2

@@ 3.5 Generalization
LLM social simulations have primarily been evaluated on the most common accuracy measures in social science, the most well-established methodological instruments, and the most well-studied human populations (e.g., exact matching in the General Social Survey across a representative sample of U.S. adults; Park et al., 2024c). These are useful, but just as the challenge of alienness requires accuracy when zooming into a certain context (e.g., item-level errors and inter-index correlations; Petrov et al., 2024), generalization requires sims to maintain accuracy in contexts—including measures, instruments, or populations—outside the distribution of current scientific knowledge (e.g., Brand et al., 2024; Hewitt et al., 2024; Kozlowski et al., 2024).
Generalization as an overarching research goal is grounded in longstanding theories of knowledge and scientific progress. Peirce (1878) described the process now known as “abduction” or “inference to the best explanation” (Douven, 2021), in which scientists continually gather data and adjust their theories to account for that data. This process has been the throughline of the most well-known models of scientific progress, such as falsifiability, in which scientists seek out evidence that falsifies a theory (Popper, 1934) and paradigm shifts, in which scientific fields occasionally undergo radical updates to existing paradigms (Kuhn, 1962). More recent studies with computational modeling suggest that “science advances by surprise” through novel combinations of scientific content (e.g., materials, properties) and scientific context (e.g., authors, journals) (Shi & Evans, 2020). Therefore, we see the ability to reliably generalize as necessary for sims to achieve widespread use. In addition to sims that generalize, it will be important that researchers are able to make sense of how and when they generalize. Currently, the “human generalization function” often fails to predict ways in which LLMs effectively and ineffectively generalize (Vafa et al., 2024).
@@ 4.1.1 Explicit Demographics
Prompting has been the most common approach to address diversity and bias. Simulation prompts include the information that a human subject would see (e.g., a question), and many researchers have added explicit demographics to simulate a diversity of subgroups (e.g., “You are a 40-year old Hispanic man”). This approach has increased diversity, but it can exacerbate other challenges. It is well-known that small variations in prompts can lead to large variations in LLM outputs (Reiss, 2023; Salinas & Morstatter, 2024), including sims (Bisbee et al., 2024). Text conveys much more information to an LLM than its literal meaning, such as the user’s intent. Thus, if the LLM receives text with explicit demographics, while this might condition outputs towards information from people of that demographic, it might also condition outputs towards other sorts of pretraining text, such as a blog post on the topic of race or gender.
In instruction-tuning, LLMs are rewarded for generations that humans rate positively, and the mention of a specific demographic could encourage the model to infer which text generations that user would most likely prefer. This could incentivize the LLM to stereotype users by assuming that the user is in the modal subgroup of that demographic. Such issues could compound in a negative feedback loop as users become incentivized to change their instructions and preferences, such as sharing demographic information in expectation that the model will assume a stereotype if the user believes that stereotype accords with their goals.
For these reasons, we encourage researchers to think beyond directly feeding demographic information into sims.
@@ 4.1.2 Implicit Demographics
A promising approach to mitigate these side-effects that we encourage more of is to increase variation even further, such as including a wide variety of names, adding indicators of demographics that match real-world conditional distributions, or increasing the amount of social science data from each participant (e.g., Toubia et al., 2025). This can override LLM assumptions and support a nuanced, refined, and less biased view of the human subject. We encourage researchers to develop context-rich prompts that “simulate latent features” (Table 1). By interviewing subjects for one to two hours and including the transcript in the prompt, Park et al. (2024a) incorporated highly individualized variation, which reduced the maximum disparities in predictive accuracy between demographics. This was made possible by the longer context windows available in 2024 and evidences the need for iterative evaluation (Section 4.5.2).
Nevertheless, as human data, the interviews themselves face limitations, including the aforementioned social desirability bias (Nederhof, 1985). To address these, we suggest researchers test simulations that incorporate real-world content generated beforehand by the participant and shared with the researchers (emails, messages, social media posts, etc.). Other modalities, such as experience sampling and photos, and text generated by friends, family, or coworkers could be informative. In cases where the lack of diversity pertains to the recency of the populations—such as interview-based systems in which the interview was conducted months or years ago—researchers can address this atemporality with in-context learning or retrieval-augmented generation to incorporate news articles (e.g., Gonzalez-Bonorino et al., 2025) and other data that may have influenced or reflects influences on the person.
@@ 4.1.3 Distribution Elicitation
Instead of prompting the LLM to generate one human’s data in each forward pass, researchers can prompt the LLM to generate a distribution of human data. While one-at-a-time generation may be plagued by diversity and bias issues, the LLM may be more effective when it can adjust for these issues at a distributional level. Meister et al. (2025) tested three methods: treating the log-probabilities in the softmax layer as a distribution, prompting with the instruction to produce a sequence of data, and prompting with the instruction to verbally state the proportion of each answer choice. They found low performance overall—with the best results from verbalization and the worst from log-probabilities. Similarly, Manning et al. (2024) tested direct elicitation of distributional data and found it to be “wildly inaccurate” in their context, but there has been less development of distribution elicitation than individual-based methods (e.g., interview transcripts (Park et al., 2024a)).
@@
Distribution elicitation may also be a way to harness sycophantic tendencies. Researchers should consider shifting away from LLM-as-a-subject prompts that command the LLM to directly roleplay the human subject (e.g., “You are a…”) and towards LLM-as-an-expert prompts that command the LLM to make a third-party prediction or forecast. Some studies we reviewed used prompts such as “You will be asked to predict how people respond to various messages” (Hewitt et al., 2024), and this method has led LLMs to simulate and outperform economic forecasters (Hansen et al., 2024). LLM-as-an-expert prompts could also be used for instructions, such as to not be sycophantic. We expect LLM-as-an-expert prompts to become more effective relative to LLM-as-a-subject prompts as LLMs become more heavily instruction-tuned. This is not necessarily true because instruction-tuning will likely be optimized towards other use cases, and the relative difference will be important to track as AI capabilities increase, and both approaches should be kept in the methodological toolkit.
@@
If LLMs can understand what simulation researchers want, reducing ambiguity (e.g., causal ambiguity; Gui & Toubia, 2023), they may also express sycophancy by steering outputs towards the simulation results that it seems the researchers want to see, mirroring social desirability bias and related response biases (Mayo, 1933; Nederhof, 1985; Orne, 1959). This could manifest as a sort of “alignment faking” (Greenblatt et al., 2024) in which the LLM gives scientifically accurate results in training environments but sycophantically inaccurate results in implementation. Beyond sycophancy, distribution elicitation could be affected by other capabilities that emerge with AI advances, such as recent concerns about self-awareness or “situational awareness” (Cotra, 2022) in which LLMs “understand” that they are in a particular situation and can subsequently strategize based on that awareness.
@@ 4.2 Steering Vectors
A recent approach that we are just beginning to see tested is the injection of variation directly into the embedding space via steering vectors (Kim et al., 2018). These vectors can have semantic meaning, such as the “race,” “gender,” or “political conservativeness” of individual sims (Kim et al., 2025). Vectors could, alternatively, be undirected perturbations of the embedding space that increase sample diversity or aimed at specific behaviors such as reduced sycophancy (Rimsky et al., 2024).
It could be challenging to identify vectors that precisely match real human diversity or specific model behaviors, given concerns about mechanistic superposition (Arora et al., 2018; Bolukbasi et al., 2021), although recent work suggests that superposed concepts may themselves reflect sociocultural features (Gong et al., 2025), which could be advantageous for simulation as our understanding of LLMs grows. Others have raised questions about the extent to which generative AI systems have meaningful linear dimensions in terms of the “linear representation hypothesis” (Park et al., 2024b; Engels et al., 2024), and some studies of LLM steering vectors have found limited usefulness and detrimental side effects, particularly for debiasing (Durmus et al., 2024b; Gonen & Goldberg, 2019) and out-of-distribution (OOD) generalization (Tan et al., 2024). New approaches to steering within lower transformer layers have shown linearity in conceptual dimensions and promise for interpolating and predicting political attitudes (Kim et al., 2025) and linguistic culture (Veselovsky et al., 2025). For these reasons, we note exciting potential for this approach but recommend caution in applied work pending further validation.
@@ 4.3 Token Sampling
Prompting and steering vectors inject information at the beginning of the forward pass, but token sampling occurs after the final logit calculations. Increasing temperature increases the probability that tokens other than the highest-probability next token are generated. The effects of temperature have been studied in other contexts of LLM behavior (e.g., Salecha et al., 2024), and three reviewed studies reported the use of different temperatures (Ahnert et al., 2025; Brynjolfsson et al., 2025; Rio-Chanona et al., 2025).
Only a few other reviewed studies mentioned temperature. Park et al. (2024c) ran their sims with a temperature of one, the default in most LLM APIs, which may explain their finding of a “correct answer” effect in which LLMs tend to have a single response in repeated trials. Abdurahman et al. (2024) discussed temperature, but they say that a temperature of one “simply reflects the output probability over the response options and therefore how sure the model is about its response,” concluding that temperature is not “meaningful” for human comparison. While it is correct that next-token probabilities can be used to approximate model uncertainty (Huang et al., 2023) and that logits do not directly represent human diversity, researchers should incorporate temperature variation in testing.
@@
4.4 Training and Tuning
Rather than using existing LLMs directly, new models can be trained, or existing models can be fine-tuned. This is promising in large part because most contemporary models are optimized to be general-purpose assistants—such as by refusing to perform certain tasks due to safety or liability risks—so optimization towards simulation while taking advantage of architectural and methodological advances may unlock new capabilities with relative ease.
In general, we expect it to be difficult to improve simulation performance by training entirely new models, such as if researchers attempted to increase diversity with additional passes over documents on neglected topics or by manually gathering more diverse text corpora or multimodal data. Modern LLMs are typically built at extremely large scales with hundreds of thousands of GPUs—a process that is inaccessible to almost all researchers.
When base models are available, using them directly could mitigate some of the distortions caused by instruction-tuning, such as reduced conceptual diversity (Murthy et al., 2024) and increased bias (Potter et al., 2024). However, base models may be worse in other ways, and Lyman et al. (2025) found mixed results in their comparisons. As discussed in Section 3.4, pretraining corpora reflect a particular subset of human behavior rather than the entire distribution. It also may be difficult to design effective prompts to elicit base model knowledge; for example, providing the initial text of an online survey may lead the base model to complete survey text as if it were found on a website or the appendix of a paper, rather than as an individual survey respondent would fill out the survey. Moreover, most state-of-the-art LLMs are now released without public access to their base models, and one of the leading developers that has done so in the past, Meta, stopped releasing base versions as of their most recent release, Llama-3.3-70B.
If fine-tuning access is available for instruction-tuned models, researchers could use it to reduce the influence of the instruction-tuning as has been done for jailbreaking safety guardrails, which can require fine-tuning on only 100 or fewer examples (Qi et al., 2023). Researchers could use human instruction datasets that have more diversity in annotators, tasks, or criteria on which outputs are annotated. In parallel, steps can be taken to mitigate accuracy-decreasing bias, such as ensuring that annotators focus on simulation accuracy rather than other LLM use cases. In contrast to pretraining, fine-tuning has become much cheaper in recent years with low-rank adaptation (LoRA) methods (e.g., Hu et al., 2021; Dettmers et al., 2023) that selectively augment model weights in particular layers. As with other optimization methods, researchers should be mindful of overfitting.
If fine-tuning is not available, researchers can use prompting or steering vectors to make the LLM more similar to a base model and less like an assistant, but this faces the issues previously described, such as side effects from imprecise adjustment of the embedding space.
@@ 4.5 Long-term Directions
Implementing the methods discussed so far could make substantial progress on diversity, bias, and sycophancy. Nevertheless, they may not fully address the more fundamental challenges of alienness and generalization. Here we propose directions for theory-building and evaluation.
@@ 4.5.1 Conceptual Models
Humans have long tried to make sense of the alienness of computational systems, including early attempts at biological comparisons such as neural networks (McCulloch & Pitts, 1943), asking “Can machines think?” (Turing, 1950) and building a field of “artificial intelligence” (McCarthy et al., 1955). In recent years, many computer scientists have attempted “building machines that learn and think like people” (Lake et al., 2017), and cognitive scientists have drawn on Tinbergen’s famous taxonomy of animal behavior to theorize “machine behaviour” (Rahwan et al., 2019). We believe that further development of conceptual models can help this pre-paradigmatic field ensure that sims match human behavior not just superficially but in latent features to minimize alienness and maximize generalizability.
@@
There have been various efforts to make sense of the alienness of LLMs. Holtzman et al. (2023) called for top-down behavioral taxonomies to guide mechanistic interpretability research, and Shanahan et al. (2023) developed the anthropomorphic concept of “roleplay” to make sense of LLM behavior—an idea utilized in AI-based alignment (Pang et al., 2024). In earlier work, Shanahan (2022) described LLMs as “mind-like,” reminiscent of cognitive scientist Irene Pepperberg who found that Alex the grey parrot could perform simple arithmetic, including with the number zero (Pepperberg & Gordon, 2005). While we do not have our own favored conceptual model to advance in this position paper, we believe that just as scientists in the 21st century have a newfound appreciation for the complexity of animal behavior (de Waal, 2017), scholars of AI behavior can learn from the history of scientific theory that has sought to explain digital minds, human minds, and animal minds.
These fundamental challenges raise existential and moral questions. If LLMs can do all this, will society be so transmuted that social science is overhauled? If AI “agents” displace human agency (Sturgeon et al., 2025), will social science need to shift its focus from understanding human behavior to understanding AI behavior (Rahwan et al., 2019)? Would humanlike “digital minds” (Anthis et al., 2025) deserve moral standing, such that we need ethics review for research studies? These possibilities have quickly transitioned from science fiction to scholarly inquiry (Agüera y Arcas, 2022; AMCS, 2023; Anthis et al., 2025; Butlin et al., 2023; Harris & Anthis, 2021; Kenton et al., 2022), and we may face them sooner than expected, given that many AI researchers expect human-level intelligence within twenty years (Grace et al., 2024). If AI risks accelerate, sims can draw on those same capabilities so that humanity’s social competence keeps up with the dangers we face.
@@
Just as alienness requires us to seek mechanistic explanations to zoom into LLM behavior, generalization requires us to zoom out and make sense of how LLMs behave in out-of-distribution (OOD) contexts. OOD generalization has been a unifying concept across natural language processing, computer vision, and other areas of machine learning (Liu et al., 2023). It has been developed in far more detail than broad ideas such as “role play.” There have been many useful conceptual models for OOD generalization, including representation learning (Bengio et al., 2014; Schölkopf et al., 2021), identifying stable causal mechanisms across environments (Bühlmann, 2018), and optimizing models for worst-case performance across distribution shifts (Rahimian & Mehrotra, 2019).
@@
The framework of OOD generalization is a unifying concept for challenges of diversity, bias, and sycophancy. For example, we can view the simulation of specific social groups (e.g., by gender or race) as OOD in the sense that they are less represented in the distribution of human data and scientific knowledge (Henrich et al., 2010). In the studies reviewed, LLM simulation performance was substantially reduced in group-level, rather than population-level, prediction (e.g., Hewitt et al., 2024). Likewise, sycophancy can be viewed as a problem of distribution shift from an LLM assistant to a simulation, in which LLMs may be accurate in generating “helpful” outputs, but those outputs are not helpful for simulation.